# WildGaussians: 3D Gaussian Splatting in the Wild

### v1

@@ Abstract
Figure 1: WildGaussians extends 3DGS [12] to scenes with appearance and illumination changes (left). It jointly optimizes DINO-based [22] uncertainty predictor to handle occlusions (right).
@@ 1 Introduction
Recently, Neural Radiance Fields (NeRFs) [1, 20, 31, 25, 32, 21, 7, 15, 24] have addressed this challenge by learning a radiance field, which combines a density field and a viewing-direction-dependent color field. These fields are rendered using volumetric rendering [10]. Despite producing highly realistic renderings, NeRFs requires evaluating numerous samples from the field per pixel to accurately approximate the volumetric integral. Gaussian Splatting (3DGS) [12, 41, 40, 42, 13, 45] has emerged as a faster alternative. 3DGS explicitly represents the scene as a set of 3D Gaussians, which enables real-time rendering via rasterization at a rendering quality comparable to NeRFs.
@@ 2 Related work
Novel View Synthesis in Dynamic Scenes. Recent methods in novel view synthesis [20, 1, 12, 41] predominantly focus on reconstructing static environments. However, dynamic components usually occur in real-world scenarios, posing challenges for these methods. One line of work tries to model both static and dynamic components from a video sequence [16, 23, 36, 37, 8, 17, 6, 38]. Nonetheless, these methods often perform suboptimally when applied to photo collection sequences [27]. In contrast, our research aligns with efforts to synthesize static components from dynamic scenes. Methods such as RobustNeRF[27] utilize Iteratively Reweighted Least Squares for outlier verification in small, controlled settings, while NeRF On-the-go [26] employs DINOv2 features [22] to predict uncertainties, allowing it to handle complex scenes with varying occlusion levels, albeit with long training times. Unlike these approaches, our method optimizes significantly faster. Moreover, we effectively handle dynamic scenarios even with changes in illumination.
Novel View Synthesis for Unstructured Photo Collections. In real-world scenes, e.g. the unstructured internet photo collections [29], challenges arise not only from dynamic occlusions like moving pedestrians and vehicles but also from varying illumination. NeRF-W [19], a pioneering work in this area, addresses these challenges with per-image transient and appearance embeddings, along with leveraging aleatoric uncertainty for transient object removal. However, the method suffers from slow training and rendering speeds. Other NeRF-based methods followed NeRF-W extending it in various ways [31, 39]. Recent concurrent works, including our own, explore the replacement of NeRF representations with 3DGS for this task. However, integrating appearance conditioning with the locally independent 3D Gaussians presents significant challenges. VastGaussian [18] applies a convolutional network to 3DGS outputs which does not transfer to large appearance changes, as shown in the Appendix. SWAG [5] addresses this by storing appearance data in an external hash-grid-based implicit field [21], while GS-W [43] utilizes CNN features for appearance conditioning on a reference image. In contrast, our method employs a simpler and more scalable strategy by embedding appearance vectors directly within each Gaussian. This design not only simplifies the architecture but also enables us to ’bake’ the trained representation back into 3DGS after appearances are fixed, enhancing both efficiency and adaptability.
@@
3 WildGaussians: Gaussian Splatting in Uncontrolled Scenes
Our approach, termed WildGaussians, is shown in Fig. 2. To allow 3DGS-based approaches handle the uncontrolled capture of scenes, we propose two key components: (1) appearance modeling enables our approach to handle the fact that the observed pixel colors not only depend on the viewpoint but also on conditions such as the capture time and the weather. Following NeRF-based approaches for reconstructing scenes from images captured under different conditions [19, 25], we train an appearance embedding per training image to model such conditions. In addition, we train an appearance embedding per Gaussian to model local effects, e.g. active illumination of parts of the scene from lamps. Both embeddings are used to transform the color stored for a Gaussian to match the color expected for a given scene appearance. To this end, we predict an affine mapping [25] in color space via an MLP. (2) uncertainty modeling allows our approach to handle occludersduring the training stage by determining which regions of a training image should be ignored.To this end, we extract DINO v2 features [22] from training images,and pass them as input to a trainable affine transformation which predicts a per-pixel uncertainty, encoding which parts of an image likely correspond to static regions and which parts show occluders. The uncertainty predictor is optimized using the cosine similarity between the DINO features extracted from training images and renderings.
@@ 3.1 Preliminaries: 3D Gaussian Splatting (3DGS)
$\Sigma^{\prime}_{i}=\big(JW\Sigma_{i}W^{T}J^{T}\big)_{1:2,1:2}\kern 5.0pt,$ (1)
@@
$\hat{C}=\sum_{i}\alpha_{i}\hat{c}_{i}(\mathbf{r})\,,\quad\quad\text{with}\quad\quad\alpha_{i}e^{\frac{1}{2}(x-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(x-\mu^{\prime}_{i})}\kern 5.0pt,$ (2)
@@ 3.2 Appearance Modeling
Following the literature on NeRFs [19, 25, 1], we use trainable per-image embeddings $\{\mathbf{e}_{j}\}_{j=1}^{N}$, where $N$ is the number of training images, to handle images with varying appearances and illuminations, such as those shown in Fig. 1. Additionally, to enable varying colors of Gaussians under different appearances, we include a trainable embedding $\mathbf{g}_{i}$ for each Gaussian $i$. We input the per-image embedding $\mathbf{e}_{j}$, per-Gaussian embedding $\mathbf{g}_{i}$, and the base color $\bar{c}_{i}$ (0-th order SH) into an MLP $f$:
$(\beta,\gamma)=f(\mathbf{e}_{j},\mathbf{g}_{i},\bar{c}_{i})\kern 5.0pt.$ (4)
@@
$\tilde{c}_{i}=\gamma\cdot\hat{c}_{i}(\mathbf{r})+\beta\kern 5.0pt.$ (5)
@@
Note that if rendering speed is important at test time and the scene only needs to be rendered under a single static condition, it is possible pre-compute the affine parameters per Gaussian and use them to update the Gaussian’s SH parameters. This essentially results in a standard 3DGS representation [12, 41] that can be rendered efficiently.
Initialization of Per-Gaussian Emeddings $\mathbf{g}_{i}$. Initializing the embedding $\mathbf{g}_{i}$ randomly could lead to a lack of locality bias, resulting in poorer generalization and training performance, as shown in supp. mat. Instead, we initialize them using Fourier features [20, 34] to enforce a locality prior. Specifically, we first center and normalize the input point cloud to the range of $[0,1]$ using the $0.97$ quantile of the $L^{\infty}$ norm. The Fourier features of a normalized point $p$ are then obtained as a concatenation of $\sin(\pi p_{k}2^{m})$ and $\cos(\pi p_{k}2^{m})$, where $k=1,2,3$ are the coordinate indices and $m=1,\ldots,4$.
@@
$\mathcal{L}_{\text{color}}=\lambda_{\text{dssim}}\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{dssim}})\|\tilde{C}-C\|_{1}\kern 5.0pt.$ (6)
@@ 3.3 Uncertainty Modeling for Dynamic Masking
To reduce the influence of transient object and occluders, e.g., moving cars or pedestrians, on the training process we learn an uncertainty model [19, 26]. NeRF On-the-go [26] showed that using features from a pre-trained feature extractor, e.g., DINO [3, 22], increases the robustness of the uncertainty predictor. However, while working well in controlled settings, the uncertainty loss function cannot handle strong appearance changes (such as these in unconstrained image collections). Therefore, we propose an alternative uncertainty loss which is more robust to appearance changes as can be seen in Figure˜3. During training, for each training image $j$, we first extract DINOv2 [22] features. Then, our uncertainty predictor is simply a trainable affine mapping applied to the DINO features, followed by the softplus activation function. Since the features are patch-wise ($14\times 14$px), we upscale the resulting uncertainty to the original size using bilinear interpolation. Finally, we clip the uncertainty to interval $0.1,\infty)$ to ensure a minimal weight is assigned to each pixel [19, 26].
@@
$\mathcal{L}_{u}=-\log\Bigg(\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\bigg({-\frac{\|\hat{C}-C\|_{2}^{2}}{2\sigma^{2}}}\bigg)\Bigg)=\frac{\|\hat{C}-C\|^{2}_{2}}{2\sigma^{2}}+\log\sigma+\frac{\log 2\pi}{2}\kern 5.0pt.$ (7)
@@
$\mathcal{L}_{\text{dino}}(\tilde{D},D)=\min\bigg(1,2-\frac{2\tilde{D}\cdot D}{\|\tilde{D}\|_{2}\|D\|_{2}}\bigg)\kern 5.0pt,$ (8)
@@
where $\sigma$ is the uncertainty prediction for the patch. We use this loss only to optimize the uncertainty predictor - single affine transformation, without letting the gradients propagate through the rendering pipeline [26]. Further, during 3DGS training, the opacity is periodically reset to a small value to prevent local minima. However, after each opacity reset, the renderings are corrupted by (temporarily) incorrect alphas. To prevent this issue from propagating into the uncertainty predictor, we hence disable the uncertainty training for a few iterations after each ocupacity reset.
Optimizing 3DGS with Uncertainty. For NeRFs, one can use the uncertainty to directly weight the training objective [19, 26]. In our experiments, we observed that this would not lead to stable training as the absolute values of gradients are used in the densification algorithm and the (randomly) large scales lead to excessive growth. The uncertainty weighting would make the setup sensitive to the correct choice of hyperparameters. Therefore, to handle this issue, we propose to convert the uncertainty scores into a (per-pixel) binary mask for the gradient scaling is at most one:
$M=\mathbbm{1}\left(\frac{1}{2\sigma^{2}}>1\right)\kern 5.0pt,$ (10)
@@
$\mathcal{L}_{\text{color-masked}}=\lambda_{\text{dssim}}M\,\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{dssim}})M\|\tilde{C}-C\|_{1}\kern 5.0pt.$ (11)
@@ 3.4 Handling Sky
For realistic renderings of a scene under different conditions, modeling the sky is important (see Fig. 1). It is unlikely that Gaussians are created in the sky when using Structure-from-Motion points as initialization. Thus, we sample points on a sphere around the 3D scene and add them to the set of points that is used to initialize the 3D Gaussians. For an even distribution of points on the sphere, we utilize the Fibonacci sphere sampling algorithm [30], which arranges points in a spiral pattern using a golden ratio-based formula. After placing these points on the sphere at a fixed radius $r_{s}$, we project them to all training cameras, removing any points not visible from at least one camera. Details are included in supp. mat.
@@ 4 Experiments
Datasets. We evaluate our WildGaussians approach on two challenging datasets. The NeRF On-the-go dataset [26] contains multiple casually captured indoor and outdoor sequences, with varying ratios of occlusions (from 5% to 30%). For evaluation, the dataset provides 6 sequences in total. Note that there are almost no illumination changes across views in this dataset. Since 3DGS [12] cannot handle radially distorted images, we train and evaluate our method and all baselines on a version of the dataset where all images were undistorted. The Photo Tourism dataset [29] consists of multiple 3D scenes of well-known monuments. Each scene has an unconstrained collection of images uploaded by users captured at different dates and times of day with different cameras and exposure levels. The average ratio of occlusions is 3.5%.
Baselines. We compare our approach against a set of baselines. On the NeRF On-the-go dataset, which contains little illumination changes, we compare to NeRF On-the-go [26] and the original 3DGS formulation [12]. On the Photo Tourism dataset [29], we compare against the most recent state-of-the-art methods for handling strong illumination changes: NeRF-W-re [19] (open source implementation) Ha-NeRF [4], K-Planes [7], RefinedFields [11], and 3DGS [12].
Metrics. We follow common practice and use PSNR, SSIM [35], and LPIPS [44] for our evaluation. We also report training times in GPU hours as well as rendering times in frames-per-second (FPS).

### v2

@@ Abstract
Figure 1: WildGaussians extends 3DGS [14] to scenes with appearance and illumination changes (left). It jointly optimizes a DINO-based [27] uncertainty predictor to handle occlusions (right).
@@ 1 Introduction
Recently, Neural Radiance Fields (NeRFs) [1, 25, 37, 30, 38, 26, 9, 17, 29] have addressed this challenge by learning a radiance field, which combines a density field and a viewing-direction-dependent color field. These fields are rendered using volumetric rendering [12]. Despite producing highly realistic renderings, NeRFs require evaluating numerous samples from the field per pixel to accurately approximate the volumetric integral. Gaussian Splatting (3DGS) [14, 50, 49, 51, 15, 54] has emerged as a faster alternative. 3DGS explicitly represents the scene as a set of 3D Gaussians, which enables real-time rendering via rasterization at a rendering quality comparable to NeRFs.
@@ 2 Related work
Novel View Synthesis in Dynamic Scenes. Recent methods in novel view synthesis [25, 1, 14, 50] predominantly focus on reconstructing static environments. However, dynamic components usually occur in real-world scenarios, posing challenges for these methods. One line of work tries to model both static and dynamic components from a video sequence [19, 28, 43, 44, 10, 21, 7, 46]. Nonetheless, these methods often perform suboptimally when applied to photo collections [32]. In contrast, our research aligns with efforts to synthesize static components from dynamic scenes. Methods such as RobustNeRF[32] utilize Iteratively Reweighted Least Squares for outlier verification in small, controlled settings, while NeRF On-the-go [31] employs DINO v2 features [27] to predict uncertainties, allowing it to handle complex scenes with varying occlusion levels, albeit with long training times. Unlike these approaches, our method optimizes significantly faster. Moreover, we effectively handle dynamic scenarios even with changes in illumination.
Novel View Synthesis for Unstructured Photo Collections. In real-world scenes, e.g. the unstructured internet photo collections [35], difficulties arise not only from dynamic occlusions like moving pedestrians and vehicles but also from varying illumination. Previously, these issues were tackled using multi-plane image (MPI) methods [20]. More recently, NeRF-W [24], a pioneering work in this area, addresses these challenges with per-image transient and appearance embeddings, along with leveraging aleatoric uncertainty for transient object removal. However, the method suffers from slow training and rendering speeds. Other NeRF-based methods followed NeRF-W extending it in various ways [37, 47]. Recent concurrent works, including our own, explore the replacement of NeRF representations with 3DGS for this task. Some methods [33, 6] address the simpler problem of training 3DGS under heavy occlusions, or only tackling appearance changes [23, 48, 8] with no occlusions. However, the main challenge is integrating appearance conditioning with the locally independent 3D Gaussians under occlusions. VastGaussian [22] applies a convolutional network to 3DGS outputs which does not transfer to large appearance changes, as shown in the Appendix. SWAG [5] and Scaffold-GS [23] address this by storing appearance data in an external hash-grid-based implicit field [26], while GS-W [52] and WE-GS [41] utilize CNN features for appearance conditioning on a reference image. In contrast, our method employs a simpler and more scalable strategy by embedding appearance vectors directly within each Gaussian. This design not only simplifies the architecture but also enables us to ’bake’ the trained representation back into 3DGS after appearances are fixed, enhancing both efficiency and adaptability. Finally, a concurrent work, Splatfacto-W [45], uses a similar appearance MLP to combine Gaussian and image embeddings to output spherical harmonics.
@@
3 Method
Our approach, termed WildGaussians, is shown in Fig. 2. To allow 3DGS-based approaches to handle the uncontrolled capture of scenes, we propose two key components: (1) appearance modeling enables our approach to handle the fact that the observed pixel colors not only depend on the viewpoint but also on conditions such as the capture time and the weather. Following NeRF-based approaches for reconstructing scenes from images captured under different conditions [24, 30], we train an appearance embedding per training image to model such conditions. In addition, we train an appearance embedding per Gaussian to model local effects, e.g., active illumination of parts of the scene from lamps. Both embeddings are used to transform the color stored for a Gaussian to match the color expected for a given scene appearance. To this end, we predict an affine mapping [30] in color space via an MLP. (2) uncertainty modeling allows our approach to handle occluders during the training stage by determining which regions of a training image should be ignored. To this end, we extract DINO v2 features [27] from training images, and pass them as input to a trainable affine transformation which predicts a per-pixel uncertainty, encoding which parts of an image likely correspond to static regions and which parts show occluders. The uncertainty predictor is optimized using the cosine similarity between the DINO features extracted from training images and renderings.
@@ 3.1 Preliminaries: 3D Gaussian Splatting (3DGS)
$\Sigma^{\prime}_{i}=\big{(}JW\Sigma_{i}W^{T}J^{T}\big{)}_{1:2,1:2}\enspace,$ (1)
@@
$\hat{C}=\sum_{i}\alpha_{i}\hat{c}_{i}(\mathbf{r})\,,\quad\quad\text{with}\quad\quad\alpha_{i}=e^{\frac{1}{2}(x-\mu^{\prime}_{i})^{T}(\Sigma^{\prime}_{i})^{-1}(x-\mu^{\prime}_{i})}\enspace,$ (2)
@@ 3.2 Appearance Modeling
Following the literature on NeRFs [24, 30, 1], we use trainable per-image embeddings $\{\mathbf{e}_{j}\}_{j=1}^{N}$, where $N$ is the number of training images, to handle images with varying appearances and illuminations, such as those shown in Fig. 1. Additionally, to enable varying colors of Gaussians under different appearances, we include a trainable embedding $\mathbf{g}_{i}$ for each Gaussian $i$. We input the per-image embedding $\mathbf{e}_{j}$, per-Gaussian embedding $\mathbf{g}_{i}$, and the base color $\bar{c}_{i}$ (00-th order SH) into an MLP $f$:
$(\beta,\gamma)=f(\mathbf{e}_{j},\mathbf{g}_{i},\bar{c}_{i})\enspace.$ (4)
@@
$\tilde{c}_{i}=\gamma\cdot\hat{c}_{i}(\mathbf{r})+\beta\enspace.$ (5)
@@
Note that if rendering speed is important at test time and the scene only needs to be rendered under a single static condition, it is possible to pre-compute the affine parameters per Gaussian and use them to update the Gaussian’s SH parameters. This essentially results in a standard 3DGS representation [14, 50] that can be rendered efficiently.
Initialization of Per-Gaussian Embeddings $\mathbf{g}_{i}$. Initializing the embedding $\mathbf{g}_{i}$ randomly could lead to a lack of locality bias, and thus poorer generalization and training performance, as shown in the supp. mat. Instead, we initialize them using Fourier features [25, 40] to enforce a locality prior: We first center and normalize the input point cloud to the range of $[0,1]$ using the $0.97$ quantile of the $L^{\infty}$ norm. The Fourier features of a normalized point $p$ are then obtained as a concatenation of $\sin(\pi p_{k}2^{m})$ and $\cos(\pi p_{k}2^{m})$, where $k=1,2,3$ are the coordinate indices and $m=1,\ldots,4$.
@@
$\mathcal{L}_{\text{color}}=\lambda_{\text{dssim}}\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{dssim}})\|\tilde{C}-C\|_{1}\enspace.$ (6)
@@ 3.3 Uncertainty Modeling for Dynamic Masking
To reduce the influence of transient objects and occluders, e.g., moving cars or pedestrians, on the training process we learn an uncertainty model [24, 31]. NeRF On-the-go [31] showed that using features from a pre-trained feature extractor, e.g., DINO [3, 27], increases the robustness of the uncertainty predictor. However, while working well in controlled settings, the uncertainty loss function cannot handle strong appearance changes (such as these in unconstrained image collections). Therefore, we propose an alternative uncertainty loss which is more robust to appearance changes as can be seen in Figure 3. During training, for each training image $j$, we first extract DINO v2 [27] features. Then, our uncertainty predictor is simply a trainable affine mapping applied to the DINO features, followed by the softplus activation function. Since the features are patch-wise ($14\times 14$px), we upscale the resulting uncertainty to the original size using bilinear interpolation. Finally, we clip the uncertainty to the interval $[0.1,\infty)$ to ensure a minimal weight is assigned to each pixel [24, 31].
@@
$\mathcal{L}_{u}=-\log\Bigg{(}\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\bigg{(}{-\frac{\|\tilde{C}-C\|_{2}^{2}}{2\sigma^{2}}}\bigg{)}\Bigg{)}=\frac{\|\tilde{C}-C\|^{2}_{2}}{2\sigma^{2}}+\log\sigma+\frac{\log 2\pi}{2}\enspace.$ (7)
@@
$\mathcal{L}_{\text{dino}}(\tilde{D},D)=\min\bigg{(}1,2-\frac{2\tilde{D}\cdot D}{\|\tilde{D}\|_{2}\|D\|_{2}}\bigg{)}\enspace,$ (8)
@@
where $\sigma$ is the uncertainty prediction for the patch. We use this loss only to optimize the uncertainty predictor (implemented as a single affine transformation) without letting the gradients propagate through the rendering pipeline [31]. Further, during 3DGS training, the opacity is periodically reset to a small value to prevent local minima. However, after each opacity reset, the renderings are corrupted by (temporarily) incorrect alphas. To prevent this issue from propagating into the uncertainty predictor, we hence disable the uncertainty training for a few iterations after each opacity reset.
Optimizing 3DGS with Uncertainty. For NeRFs, one can use the uncertainty to directly weight the training objective [24, 31]. In our experiments, we observed that this would not lead to stable training as the absolute values of gradients are used in the densification algorithm and the excessively large magnitudes of absolute gradients lead to excessive growth. The uncertainty weighting would make the setup sensitive to the correct choice of hyperparameters. Therefore, to handle this issue, we propose to convert the uncertainty scores into a (per-pixel) binary mask such that the gradient scaling is at most one:
$M=\mathbbm{1}\left(\frac{1}{2\sigma^{2}}>1\right)\enspace,$ (10)
@@
$\mathcal{L}_{\text{color-masked}}=\lambda_{\text{dssim}}M\,\text{DSSIM}(\hat{C},C)+(1-\lambda_{\text{dssim}})M\|\tilde{C}-C\|_{1}\enspace.$ (11)
@@ 3.4 Handling Sky
For realistic renderings of a scene under different conditions, modeling the sky is important (see Fig. 1). It is unlikely that Gaussians are created in the sky when using Structure-from-Motion points as initialization. Thus, we sample points on a sphere around the 3D scene and add them to the set of points that is used to initialize the 3D Gaussians. For an even distribution of points on the sphere, we utilize the Fibonacci sphere sampling algorithm [36], which arranges points in a spiral pattern using a golden ratio-based formula. After placing these points on the sphere at a fixed radius $r_{s}$, we project them to all training cameras, removing any points not visible from at least one camera. Details are included in the supp. mat.
@@ 4 Experiments
Methods were trained and evaluated on NVIDIA A100, while the rest used NVIDIA GTX 4090.
@@
Evaluated using NeRF-W evaluation protocol [24, 18]; Source code not available, numbers from the paper, unknown GPU used.
@@
Datasets. We evaluate our WildGaussians approach on two challenging datasets. The NeRF On-the-go dataset [31] contains multiple casually captured indoor and outdoor sequences, with varying ratios of occlusions (from 5% to 30%). For evaluation, the dataset provides 6 sequences in total. Note that there are almost no illumination changes across views in this dataset. Since 3DGS [14] cannot handle radially distorted images, we train and evaluate our method and all baselines on a version of the dataset where all images were undistorted. The Photo Tourism dataset [35] consists of multiple 3D scenes of well-known monuments. Each scene has an unconstrained collection of images uploaded by users captured at different dates and times of day with different cameras and exposure levels. In our experiments we use the Brandenburg Gate, Sacre Coeur, and Trevi Fountain scenes, which have an average ratio of occlusions of 3.5%. Note, that for each dataset (both NeRF On-the-go and Photo Tourism), the test set was carefully chosen not to have any occluders.
Baselines. We compare our approach against a set of baselines. We use NerfBaselines [18] as our evaluation framework, providing a unified interface to the original released source codes while ensuring fair evaluation. On the NeRF On-the-go dataset, which contains little illumination changes, we compare to NeRF On-the-go [31], the original 3DGS formulation [14], Mip-Splatting [50], and Gaussian Opacity Fields [51]. On the Photo Tourism dataset [35], we compare against the most recent state-of-the-art methods for handling strong illumination changes: NeRF-W-re [24] (open source implementation), Ha-NeRF [4], K-Planes [9], RefinedFields [13], 3DGS [14], and concurrent works GS-W [52] and SWAG [5]. We evaluate GS-W [52] using the NeRF-W evaluation protocol [24] (see below). Our GS-W numbers thus differ from those in [52] (which conditioned on full test images).
Metrics. We follow common practice and use PSNR, SSIM [42], and LPIPS [53] for our evaluation. For the Photo Tourism dataset [35], we use the evaluation protocol proposed in NeRF-W [24, 18], where the image appearance embedding is optimized on the left half of the image. The metrics are then computed on the right half. For the NeRF On-the-go dataset [31], there is no test-time optimization. We also report training times in GPU hours as well as rendering times in frames-per-second (FPS), computed on an NVIDIA RTX 4090 unless stated otherwise.